OpenStudy (anonymous):

1/9 x= 3...kinda forgot how to do it..oops. 1/9 x= 3...kinda forgot how to do it..oops. @Mathematics

OpenStudy (anonymous):

Multiply both sides by 9 to get 9(1/9x)=9(3), and 1x is simply x. So you get x=27.

OpenStudy (anonymous):

multiply by the reciprocal: 1/9 x= 3 times both sides by 9: 9/9 x = 27 (9/9 = 1) x = 27